OpenStudy (anonymous):

Use substitution to solve the following systems of equations 9x+4y=2 2x+y=1

OpenStudy (apoorvk):

Great okay.. can you do something for me - just multiply every term of both the LHS and the RHS of the second equation by 4? What do you get?

OpenStudy (anonymous):

LHS RHS? whats that

OpenStudy (lgbasallote):

use \(\small \textbf{*substitution*}\) gotta love that word hehe

OpenStudy (anonymous):

x = -2, y = 5 GOGOGO :P

OpenStudy (anonymous):

lhs = left hand side rhs = right hand side

OpenStudy (anonymous):

how do you get that though i want to understand

OpenStudy (lgbasallote):

look at your equations: 1) 9x + 4y = 2 2) 2x + y = 1 which one has no coefficient?

OpenStudy (anonymous):

the second one

OpenStudy (lgbasallote):

which one in the second?

OpenStudy (anonymous):

y

OpenStudy (lgbasallote):

right. so isolate y

OpenStudy (lgbasallote):

what do you get if you isolate y in 2x + y = 1

OpenStudy (anonymous):

1?

OpenStudy (lgbasallote):

what happened to 2x?

OpenStudy (anonymous):

so its 2x+-2??

OpenStudy (anonymous):

idk to be honest im lost

OpenStudy (lgbasallote):

no..

OpenStudy (lgbasallote):

look at 2x + y = 1 what will you do so that only y is on the left side?

OpenStudy (anonymous):

subtract 2 from 2x?

OpenStudy (lgbasallote):

nope... let's try this 3 + y = 1 what will you do to isolate y?

OpenStudy (anonymous):

we subtract 3 from both sides?

OpenStudy (lgbasallote):

correct

OpenStudy (lgbasallote):

so how do you isolate y in 2x + y = 1

OpenStudy (anonymous):

we subtract 2x from both sides

OpenStudy (lgbasallote):

right..what does that give you

OpenStudy (anonymous):

so it will give me y=-1

OpenStudy (lgbasallote):

no you subtracted 2x from BOTH SIDES right?

OpenStudy (anonymous):

yea 2x+y=1 -2 cancels out

OpenStudy (lgbasallote):

2x - 2 does not cancel out

OpenStudy (anonymous):

its gives u -4

OpenStudy (lgbasallote):

no.. you cant combine 2x and 2

OpenStudy (lgbasallote):

remember what you said earlier..we subtract 2x from both sides...not 2

OpenStudy (anonymous):

2x+y=1 -2x -2x

OpenStudy (anonymous):

so then it cancels out

OpenStudy (anonymous):

hep

OpenStudy (turingtest):

-2x and 2x cancel on the left side, but what makes you think -2x cancels with anything on the right side?

OpenStudy (anonymous):

no im saying it cancels out on that side

OpenStudy (turingtest):

which side?

OpenStudy (anonymous):

on the left

OpenStudy (anonymous):

then its 1-2x right which equals

OpenStudy (turingtest):

which equal what...?

OpenStudy (anonymous):

-1

OpenStudy (turingtest):

why is it equal to -1 ??? no... what you have for the RHS is correct but on the left, what is left after the 2x and the -2x cancel?

OpenStudy (anonymous):

y=-1

OpenStudy (anonymous):

ugh this sucks

OpenStudy (turingtest):

you need to think of each side separately

OpenStudy (anonymous):